find the the value of k if the straight line 6x-10y+3=0 and kx-5y+8=0 are parallel
 

Grade:11

3 Answers

Khimraj
6 years ago
For parallel lines does should be equal so 6/10 = k/5 so k=3.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
Vaibhav
6 years ago
For parallel line a¹/a² =b¹/b²is unequal to c¹/c². Here a¹ = 6, a² = k, b¹ = -10, b² = - 5.So, on putting the values in the above equation, we get k = 3. If you like my answer then please approve it.
Sriesh
6 years ago
When 2 lines are parallel their slope are equal so for line 6x-10y +3 the slope is 6/10(by the expression y=mx+c)Now slope of the other line is K/5As the lines are parallel both the slopes are equalThus 6/10=k/5K=3